OpenStudy (anonymous):

The Richter scale converts seismographic readings into numbers for measureing trhe magnitude of an earthquake according to this function M(x)=log(x/xO)...What is the magnitude of an earthquake whose seismographic reading is 0.94 millimeters at a distance of 100 kilometers from its epicenter? round the answer to four decimal places? How do I set up this Logarthmic problem???? HELP HELP

OpenStudy (anonymous):

I want to mainly see how I do the problem it's not so much the answer but how the answer is achieved if you can I would appreciate it?

OpenStudy (anonymous):

what is x? what is \[x_0\]?

OpenStudy (anonymous):

is x the seismographic reading? and what does 100 km have to do with it? i thought \[x_0\] was a constant. need a little more info on this to do the problem

OpenStudy (anonymous):

oh sorry x0=10^-3

OpenStudy (anonymous):

ok and what about the 100 km?

OpenStudy (anonymous):

LOL..so sorry just a tad bit of very important infor I left out

OpenStudy (anonymous):

uh that's all it says about the 100 k I know it's a log question but Im not sure what would M(x) be

OpenStudy (anonymous):

first off \[\log(\frac{x}{10^{-3}})=\log(x)-(-3)=\log(x)+3\]

OpenStudy (anonymous):

so it looks like you should just take \[M(x)=\log(.94)+3\] but i am don't know about this 100 km thing so i am going to get off ad let some one else try

OpenStudy (anonymous):

k thank you it is a doosey